Circular No. 5813

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

6 CEPHEI L. Szabados and M. Kun, Konkoly Observatory, Budapest, communi-

cate: “The well-known Be star 6 Cep = HR 8171 = HD 203467 has been remarkably stable from a photometric point of view for years, the yearly mean values being as follows: 1989, V = 5.20, B–V = –0.06, U–B = –0.58; 1990, 5.22, –0.05, –0.56; 1991, 5.23, –0.06, –0.54; 1992, 5.22, –0.05, –0.54. In 1993, however, a definite brightening is observed, which may indicate the commencement of a new activity phase. The recently obtained data are: 1993 May 29.86 UT, V = 5.10, B–V = –0.01, U–B = –0.61; June 9.84, 5.08, +0.01, –0.62; 9.94, 5.09, –0.01, –0.62. These magnitudes have been measured at Konkoly using 7 Cep = HR 8227 as the comparison star, constancy of which was regularly checked against HR 8133. Spectroscopic observations are recommended in order to study the mass motion in the stellar atmosphere while this episode lasts.”

CYGNUS X-1 B. A. Harmon, G. J. Fishman, and C. A. Wilson, Marshall Space Flight

Center, NASA; W. S. Paciesas, University of Alabama, Huntsville; and J. C. Ling and W. A. Wheaton, Jet Propulsion Laboratory, California Institute of Technology, report for the Compton Observatory/BATSE Team: “Hard x-ray emission from the black-hole candidate Cygnus X-1, measured using the Earth occultation technique, has decreased gradually from a flux level of 0.12 to ∼ 0.08 photon cm^{−2} s^{−1} (45- to 140-keV band) from early May to ∼ May 23. As of June 8, the flux remained at 0.07-0.09 photon cm^{−2} s^{−1}. The two flux levels correspond roughly to the γ_{2} (higher flux) and γ_{1} (lower flux) states observed in 1979-1980 by HEAO-3 (Ling et al. 1987, Ap.J. 321, L117). These observations indicate the possibility of a γ_{2} to γ_{1} transition which has not been seen previously.”

NOVA AQUILAE 1993 Further CCD photometry from H. Mikuž, Ljubljana, Slovenia (cf.

IAUC 5807): May 31.986 UT, V = 8.54 ± 0.02, B–V = +0.50 ± 0.04; June 2.063, V = 8.76 ± 0.02; 6.020, V = 9.20 ± 0.04 (strong moonlight).

1993 June 11 Daniel W. E. Green